OpenStudy (anonymous):

**GIVING MEDAL** What is f(-5) if f(x) = |2x - 1| + 10? 21 19 -1 -21

OpenStudy (anonymous):

f is the name of the recipe x is the ingredient |x + 1| is the set of instructions (what to do with the ingredient) In this case, it says: take the ingredient (x), add 1 to it (x+1), then remove the sign (that is what the vertical bars say: it is called an "absolute" sign, and it tells you to drop the sign of whatever is inside). | -7 | = 7 | + 14 | = 14 | square root of 4 | = 2 (without the bars, the square root of 4 COULD be -2, because (-2)(-2) = 4, but with the vertical bars, you are forced to use only the non-negative value) --- f(-5) is a shorthand way of saying: Use the recipe called "f" and use "-5" as the ingredient: Take the ingredient: -5 add 1 to it: -5 + 1 = -4 remove the sign |-4| = 4

OpenStudy (anonymous):

@angeleye I understand what you're saying, but is 4 the outcome or is it the beginning of finding the answer?

OpenStudy (anonymous):

@austinL @JjJustin please help if you can!

OpenStudy (***[isuru]***):

hi,," just use this... the basic meaning of |2x - 1| is.... u will use -(2x -1) instead of |2x - 1| if "x is lesser than 1/2 and u will use (2x -1) instead of |2x - 1| if "x" is larger than or equal to 1/2 so........... when u need to find f(-5) u need to substitute (-5) instead of "x" in f(x) but when x = (-5) in f(x) x become lesser than 1/2 so... u will have to change f(x) in to.. |2x - 1| + 10 ----> -(2x -1) + 10 according to the rule above now u can substitute (-5) instead of "x" in -(2x -1) + 10 which is.. -{ [2*(-5)] - 1} +10 now ... wt is ur answer bro ?

OpenStudy (anonymous):

@***[ISURU]*** My answer then would be 19?

OpenStudy (anonymous):

c'mon its easy just plug the values after putting x=-5 what we get? |2(-5)-1|+10 |-10-1|+10 |-11|+10=21 where are you having problem?

OpenStudy (***[isuru]***):

nope... -{ [2*(-5)] - 1} +10 - ( -11) +10 11 + 10 wt is ur answer now ?

OpenStudy (austinl):

\(f(x)=|2x-1|+10\) \(f(-5)\) means that we will merely input "-5" for all instances of "x" in the function. \(f(-5)=|2(-5)-1|+10\Rightarrow~|-10-1|+10\) \(|-10-1|=|-11|\Rightarrow~11\) So therefore we have, \(f(-5)=11+10+~?\)

OpenStudy (austinl):

\(\color{red}{\text{Correction}}\) \(f(-5)=11+10=~?\)